OpenStudy (anonymous):

Use the Rational Zeros Theorem to write a list of all possible rational zeros of the function. f(x) = -2x4 + 4x3 + 3x2 + 18

OpenStudy (anonymous):

Try this site for an explanation and example: http://www.cliffsnotes.com/study_guide/Rational-Zero-Theorem.topicArticleId-257309,articleId-257204.html

OpenStudy (campbell_st):

ok... you need to find the factors of 18 called p \[1 \times 18, -1 \times -18, 2 \times 9, -2 \times -9, 3 \times 6, -3 \times - 6\] and factors of -2 called q \[-1 \times 2, 1 \times -2 \] then find all the combinations of p/q then find f(p/q) if there is a zero remainder p/q is a factor of f(x)

OpenStudy (anonymous):

im confused lol.. this type of math ive never really understood?

OpenStudy (anonymous):

so the answer would be 1,2,-1,-2?